# LAPTH-015/24 Spherical symmetric solutions of

### v1

@@
Gérard Clémenta and Khireddine Nouicerb
a LAPTh, Université Savoie Mont Blanc, CNRS, F-74940 Annecy, France
b LPTh, Department of Physics, University of Jijel,
@@ 1 Introduction
In the pursuit of understanding the universe’s accelerated expansion [2, 3], numerous alternatives have been proposed, both within the framework of general relativity and beyond. One such alternative involves introducing a new matter component termed dark energy, characterized by negative pressure. While the cosmological constant $\Lambda$, originally introduced by Einstein, remains the primary candidate for dark energy due to its alignment with observational data, challenges such as the fine-tuning and coincidence problems persist [4, 5, 6, 7]. On the other hand, a different path follows the idea that general relativity fails to describe the universe at larger scales and hence it should be modified. These modifications often propose alterations to fundamental principles of physics, seeking to reconcile observational evidence with theoretical predictions [8, 9]. Consequently, alternative models like scalar tensor theories [10, 11], Einstein-æther theory [12, 13], and higher derivative theories of gravity like $f(R)$ theories [14] have been explored. Other kinds of modified theories are the modified teleparallel equivalent of general relativity, $f(T)$ gravity where the Ricci scalar is replaced by the torsion scalar, and the theory grounded in Weyl geometry, termed $f(Q)$ gravity [15]. The latter describes gravitational effects in terms of non-metricity, providing a broader framework than traditional Riemannian geometry. These theories offer different perspectives on gravity that may lead to new insights into cosmological and astrophysical phenomena.
@@ 2 Static spherically symmetric solutions with Maxwell source
$y(r)=\frac{q^{2}\mu^{2}}{4r^{4}}-\frac{\mu^{2}}{2r^{2}}+k_{1}+k_{2}\ln\left|\*dfrac{r}\mu\right|-\frac{\Lambda r^{2}}{3}.$ (2.19)
@@ 3.2.1 Subcase 2a
with $k_{1}>-1$ for $y(\infty)>0$. For $-1<k_{1}<0$, the spacetime [2.13) is a black hole, with a single horizon shielding the singularity $r=0$. For $k_{1}>0$ the singularity is naked. For $k_{1}=0$, the singularity-free metric
@@ 5 Time-dependent vacuum solutions
(where ˙$\dot{}=d/dt$), Harada showed that the equations of conformal Killing gravity may be first integrated to the second order equation
@@
If $\alpha=0$ we recover the usual Friedmann equation. Thus, the integration constant $\alpha$ measures the deviation of conformal Killing gravity from cosmological general relativity.
@@ 5.1 $\alpha>0$
If $\Lambda<-6\sqrt{-k\alpha}$, the two roots are real, so that the evolution is time symmetrical, with either the universe contracting to a minimum radius at $a=a_{+}$ and then expanding again to infinity in a finite time interval $2t_{0}$; or blowing up from a big bang at $t=-t_{0}$ to a maximum radius at $a=a_{-}$ and then recontracting to a big crunch at $t=t_{0}$. *

### v2

@@
[…]
${}^{a}$ LAPTh, Université Savoie Mont Blanc, CNRS, F-74940 Annecy, France
${}^{b}$ LPTh, Department of Physics, University of Jijel,
@@ 1 Introduction
In the pursuit of understanding the universe’s accelerated expansion [2, 3], numerous alternatives have been proposed, both within the framework of general relativity and beyond. One such alternative involves introducing a new matter component termed dark energy, characterized by negative pressure. While the cosmological constant $\Lambda$, originally introduced by Einstein, remains the primary candidate for dark energy due to its alignment with observational data, challenges such as the fine-tuning and coincidence problems persist [4, 5, 6, 7]. On the other hand, a different path follows the idea that general relativity fails to describe the universe at larger scales and hence it should be modified. These modifications often propose alterations to fundamental principles of physics, seeking to reconcile observational evidence with theoretical predictions [8, 9]. Consequently, alternative models like scalar tensor theories [10, 11], Einstein-æther theory [12, 13], and higher derivative theories of gravity like $f(R)$ theories [14, 15] have been explored. Other kinds of modified theories are the modified teleparallel equivalent of general relativity, $f(T)$ gravity where the Ricci scalar is replaced by the torsion scalar, and the theory grounded in Weyl geometry, termed $f(Q)$ gravity [16]. The latter describes gravitational effects in terms of non-metricity, providing a broader framework than traditional Riemannian geometry. These theories offer different perspectives on gravity that may lead to new insights into cosmological and astrophysical phenomena.
@@ 2 Static spherically symmetric solutions with Maxwell source
It was shown in [18] that the Harada field equations (2.1) are equivalent to the Einstein equations
$R_{\mu\nu}-\displaystyle\frac{1}{2}Rg_{\mu\nu}=8\pi G(T_{\mu\nu}+K_{\mu\nu})$ (2.6)
modified by the addition to the matter energy-momentum tensor of a divergence- free ($\nabla_{\nu}K^{\mu\nu}=0$) conformal Killing tensor $K_{\mu\nu}$ of trace $K$ satisfying the equation
$\nabla_{\mu}K_{\nu\rho}+\nabla_{\nu}K_{\rho\mu}+\nabla_{\rho}K_{\mu\nu}-\frac{1}{6}\left(g_{\nu\rho}\partial_{\mu}+g_{\rho\mu}\partial_{\nu}+g_{\mu\nu}\partial_{\rho}\right)K=0.$ (2.7)
This formulation will not be used in the following.
@@
$y(r)=\frac{q^{2}\mu^{2}}{4r^{4}}-\frac{\mu^{2}}{2r^{2}}+k_{1}+k_{2}\ln\left|\displaystyle\frac{r}{\mu}\right|-\frac{\Lambda r^{2}}{3}.$ (2.21)
@@ 3.2.1 Subcase 2a
with $k_{1}>-1$ for $y(\infty)>0$. For $-1<k_{1}<0$, the spacetime [2.15) is a black hole, with a single horizon shielding the singularity $r=0$. For $k_{1}>0$ the singularity is naked. For $k_{1}=0$, the singularity-free metric
@@ 5 Time-dependent vacuum solutions
(where ˙$\dot{}=d/dt$), Harada [17] showed that the equations of conformal Killing gravity may be first integrated to the second order equation
@@
This equation was previously obtained through an alternative method in [18] (our constant $\alpha$ is equivalent to their constant $-C/6$) . If $\alpha=0$ we recover the usual Friedmann equation. Thus, the integration constant $\alpha$ measures the deviation of conformal Killing gravity from cosmological general relativity [18].
@@ 5.1 $\alpha>0$
If $\Lambda<-6\sqrt{-k\alpha}$, the two roots are real, so that the evolution is time symmetrical, with either the universe contracting to a minimum radius at $a=a_{+}$ and then expanding again to infinity in a finite time interval $2t_{0}$; or blowing up from a big bang at $t=-t_{0}$ to a maximum radius at $a=a_{-}$ and then recontracting to a big crunch at $t=t_{0}$.